(Exa) 1.Find what is the probability of value of X being <= 110 for the Normal Distribution defined by average of 92, and standard deviation of 14.

2.What would be the probability of X being > 110. mean is 92 and st. deviation is 14 (Remember total probability is 1!!)
1
Expert's answer
2020-12-08T06:33:24-0500

1.

"P(\le110)=P(Z\le \frac{110-92}{14})"

"=P(Z\le1.29)=0.9015" from z tables

"P(\le110)=0.9015"

2.

"P(>110)=1-P(\le110)"

"=1-0.9015=0.0985"